Solve x for the following 4*3^(x-1) = 5^(2x+3)?
I believe this breaks down to the following
log(4) + (x-1)log(3) = 2xlog(5) + log(3)

I have been trying to do this equation by working the individual logs out and then adding/subtracting as needed but I cannot get the right answer.

Answer by ankor@dixie-net.com(22740)

Solve x for the following 4 * 3^(x-1) = 5^(2x+3)?
I believe this breaks down to the following
log(4) + (x-1)log(3) = 2xlog(5) + log(3)
:
Wouldn't it be
log(4) + (x-1)log(3) = (2x+3)log(5)
: